OpenStudy (anonymous):

Let cos data = -3/5, and data>0 is in quadrant three. find the exact values of the remaining five trigonometric functions. **please explain how you got your answer**

OpenStudy (anonymous):

\[\cos \theta=-\frac{ 3 }{ 5 } and \theta >0\]